OpenStudy (anonymous):

find an equation of the line passing through the point (-2,4) and parallel to the line given by the equation 2x-3y=11, write your answer in slope intercept form.

OpenStudy (anonymous):

parallel lines have the same slope, so if we re-write your given equation in form: y=2/3 x -11/3 your slope will be 2/3 right?

OpenStudy (anonymous):

now, to pass through he point (-2,4) means: y-4=2/3 (x-(-2))

OpenStudy (anonymous):

Their initial equation was 2x-3y=11

OpenStudy (anonymous):

or y=2x/3 +16/3 see it?

OpenStudy (anonymous):

I re-wrote give eq as follow (slope-intercept form): -3y=11-2x y=2/3 x -11/3

OpenStudy (anonymous):

Oh yeah sorry me just being stupid thanks

OpenStudy (anonymous):

no worries... :)

OpenStudy (anonymous):

thank you :)